OpenStudy (anonymous):

Solve the system by the substitution method. xy = 12 x2 + y2 = 40

OpenStudy (anonymous):

\[x^2+y^2=40\]\[x*y=12\] so \[(x-y)^2=16\]\[(x+y)^2=64\]\[x-y=\pm4\] and, \[x+y=\pm8\]

OpenStudy (anonymous):

so by solving the above equations that are x+y=8 and, x-y=8 we get the following ordered pairs (x,y)=\[(6,2), (-6,-2)\]